40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15768     Entry time: Fri Jan 15 17:04:45 2021
Author: gautam 
Type: Update 
Category: LSC 
Subject: Messed up LSC sensing 

I want to lock the PRFPMI again (to commission AS WFS). Have had some success - but in doing characterization, I find that the REFL port sensing is completely messed up compared to what I had before. Specifically, MICH and PRCL DoFs have no separation in either the 1f or 3f photodiodes. 

  • A sensing line driven in PRCL doesn't show up in the AS55 photodiode signal - this is good and as expected.
  • For MICH - I set the MICH--->PRM actuation matrix element so as to minimize the height of the peak at the MICH drive frequency that shows up at the PRCL error point. My memory is that I used to be able to pretty much null this signal in the past, but I can't find a DTT spectrum in the elog as evidence. Anyways, the best effort nulling I can achieve now still results in a large peak at the PRCL error point. Since the sensing matrix doesn't actually make any sense, idk if it is meaningful to even try and calibrate the above qualitative statement into quantitative numbers of cross coupling in meters.
  • With the PRMI locked on 1f error signals (ETMs misaligned, PRCL sensed with REFL11_I, MICH sensed with AS55_Q) - I tried tweaking the digital demod phase of the REFL33 and REFL165 signals. But I find that the MICH and PRCL peaks move in unison as I tweak the demod phase. This suggests to me that both signals are arriving optically in phase at the photodiode, which is weird.
  • The phenomenon is seen also in the REFL11 signal.

I did make considerable changes to the RF source box, and so now the relative phase between the 11 MHz and 55 MHz signals is changed compared to what it was before. But do we really expect any effect even in the 1f signal? I am not able to reproduce this effect in simulation (Finesse), though I'm using a simplified model. I attach two sensing matrices to illustrate what i mean:

  1. Attachment #1 is in the PRFPMI state, with the IFO on RF control (CARM on REFL11, PRCL on REFL165_I, MICH on REFL165_Q, DARM on AS55_Q). 
  2. Attachment #2 is between the transition to RF control (CARM and DARM on ALS, PRCL on REFL165_I, MICH on REFL165_Q). The CARM offset is ~4nm (c.f. the linewidth of ~20pm), so the circulating power in the arm cavities is low.
Attachment 1: PRFPMI_Jan12sensMat.pdf  45 kB  | Hide | Hide all
PRFPMI_Jan12sensMat.pdf
Attachment 2: PRMI3f_ALS_Jan11_largeOffsetsensMat.pdf  38 kB  | Hide | Hide all
PRMI3f_ALS_Jan11_largeOffsetsensMat.pdf
ELOG V3.1.3-